OpenStudy (anonymous):

Write the standard form of the line that contains a slope of 1/2 and y-intercept of 3??

OpenStudy (lochana):

well, The standard for of any line id Ay + Bx + C = 0 if you draw a graph out of all given information. you will find the equation of that graph is y = (1/2)x + 3 now you can make it in standard form 2y = x + 6 2y - x - 6 = 0

OpenStudy (anonymous):

So that's it? I don't have to do anything else with it??

OpenStudy (lochana):

No. that's it. :)

OpenStudy (anonymous):

Oh, well thank you.